### v2

# GVE-Louvain: Fast Louvain Algorithm for Community Detection in Shared Memory Setting

Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents an optimized parallel implementation of Louvain, a high quality community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Louvain, which we term as GVE-Louvain, outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively - achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Louvain improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
Community detection, Parallel Louvain algorithm
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Community detection is the problem of uncovering the underlying structure of complex networks, i.e., identifying groups of vertices that exhibit dense internal connections but sparse connections with the rest of the network, in an unsupervised manner. It is an NP-hard problem with numerous applications in domains such as drug discovery, protein annotation, topic discovery, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). The Louvain method (Blondel et al., 2008) is a popular heuristic-based approach for community detection, with the modularity metric (Newman, 2006) being used to measure the quality of communities identified.
In recent years, the collection of data and the relationships among them, represented as graphs, have reached unmatched levels. This has necessitated the design of efficient parallel algorithms for community detection on large networks. The multicore/shared memory setting is crucial for community detection due to its energy efficiency and the prevalence of hardware with extensive DRAM sizes. However, many of the current algorithms for community detection are challenging to parallelize due to their irregular and inherently sequential nature (Halappanavar et al., 2017), in addition to the complexities of handling concurrency, optimizing data access, reducing contention, minimizing load imbalance. Existing studies on Louvain propose algorithmic optimizations and parallelization techniques, but do not address optimization for the aggregation phase of the Louvain algorithm, which emerges as a bottleneck after the local-moving phase of the algorithm has been optimized. Moreover, these optimization techniques are scattered over a number of papers, making it difficult for a reader to get a grip over them.
1.1. Our Contributions
This report introduces GVE-Louvain, an optimized parallel implementation of Louvain [^1] https://github.com/puzzlef/louvain-communities-openmp for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-Louvain outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively, achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-Louvain exhibits an average performance scaling of 1.6×1.6$\times$1.6 ×.
2. Related work
The Louvain method is a greedy modularity-optimization based community detection algorithm, and is introduced by Blondel et al. from the University of Louvain (Blondel et al., 2008). It identifies communities with resulting high modularity, and is thus widely favored (Lancichinetti and Fortunato, 2009). Algorithmic improvements proposed for the original algorithm include early pruning of non-promising candidates (leaf vertices) (Ryu and Kim, 2016; Halappanavar et al., 2017; Zhang et al., 2021; You et al., 2022), attempting local move only on likely vertices (Ryu and Kim, 2016; Ozaki et al., 2016; Zhang et al., 2021; Shi et al., 2021), ordering of vertices based on node importance (Aldabobi et al., 2022), moving nodes to a random neighbor community (Traag, 2015), threshold scaling (Lu et al., 2015; Naim et al., 2017; Halappanavar et al., 2017), threshold cycling (Ghosh et al., 2018b), subnetwork refinement (Waltman and Eck, 2013; Traag et al., 2019), multilevel refinement (Rotta and Noack, 2011; Gach and Hao, 2014; Shi et al., 2021), and early termination (Ghosh et al., 2018b).
To parallelize the Louvain algorithm, a number of strategies have been attempted. These include using heuristics to break the sequential barrier (Lu et al., 2015), ordering vertices via graph coloring (Halappanavar et al., 2017), performing iterations asynchronously (Que et al., 2015; Shi et al., 2021), using adaptive parallel thread assignment (Fazlali et al., 2017; Naim et al., 2017; Sattar and Arifuzzaman, 2019; Mohammadi et al., 2020), parallelizing the costly first iteration (Wickramaarachchi et al., 2014), using vector based hashtables (Halappanavar et al., 2017), and using sort-reduce instead of hashing (Cheong et al., 2013). Platforms used range from an AMD multicore system (Fazlali et al., 2017), and Intel’s Knight’s Landing, Haswell (Gheibi et al., 2020), SkylakeX, and Cascade Lake (Hossain and Saule, 2021). Other approaches include the use of MapReduce in a BigData batch processing framework (Zeitz, 2017). It should however be noted though that community detection methods such as the Louvain that rely on modularity maximization are known to suffer from resolution limit problem. This prevents identification of communities of certain sizes (Ghosh et al., 2019).
A few open source implementations and software packages have been developed for community detection. Vite (Ghosh et al., 2018a) is a distributed memory parallel implementation of the Louvain method that incorporates several heuristics to enhance performance while maintaining solution quality, while Grappolo (Halappanavar et al., 2017) is a shared memory parallel implementation. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes a parallel implementation of the Louvain algorithm.
3. Preliminaries
Consider an undirected graph $G(V,E,w)$ with $V$ representing the set of vertices, $E$ the set of edges, and $w_{ij}=w_{ji}$ denoting the weight associated with each edge. In the case of an unweighted graph, we assume unit weight for each edge ($w_{ij}=1$). Additionally, the neighbors of a vertex $i$ are denoted as $J_{i}=\{j\ |\ (i,j)\in E\}$, the weighted degree of each vertex as $K_{i}=\sum_{j\in J_{i}}w_{ij}$, the total number of vertices as $N=|V|$, the total number of edges as $M=|E|$, and the sum of edge weights in the undirected graph as $m=\sum_{i,j\in V}w_{ij}/2$.
3.1. Community detection
Disjoint community detection is the process of identifying a community membership mapping, $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is assigned a community-id $c\in\Gamma$, where $\Gamma$ is the set of community-ids. We denote the vertices of a community $c\in\Gamma$ as $V_{c}$, and the community that a vertex $i$ belongs to as $C_{i}$. Further, we denote the neighbors of vertex $i$ belonging to a community $c$ as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, the sum of those edge weights as $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$, the sum of weights of edges within a community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of a community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec, 2021).
3.2. Modularity
Modularity serves as a metric for assessing the quality of communities identified by heuristic-based community detection algorithms. It is computed as the difference between the fraction of edges within communities and the expected fraction if edges were randomly distributed, yielding a range of $[-0.5,1]$ where higher values indicate better results (Brandes et al., 2007). The modularity $Q$ of identified communities is determined using Equation 1, where $\delta$ represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise). The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, can be computed using Equation 2.
(1) $Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
(2) $\Delta Q_{i:d\rightarrow c}=\frac{1}{m}(K_{i\rightarrow c}-K_{i\rightarrow d})-\frac{K_{i}}{2m^{2}}(K_{i}+\Sigma_{c}-\Sigma_{d})$
3.3. Louvain algorithm
The Louvain method (Blondel et al., 2008) is a modularity optimization based agglomerative algorithm for identifying high quality disjoint communities in large networks. It has a time complexity of $O(L|E|)$ (with $L$ being the total number of iterations performed), and a space complexity of $O(|V|+|E|)$ (Lancichinetti and Fortunato, 2009). The algorithm consists of two phases: the local-moving phase, where each vertex $i$ greedily decides to move to the community of one of its neighbors $j\in J_{i}$ that gives the greatest increase in modularity $\Delta Q_{i:C_{i}\rightarrow C_{j}}$ (using Equation 2), and the aggregation phase, where all the vertices in a community are collapsed into a single super-vertex. These two phases make up one pass, which repeats until there is no further increase in modularity (Blondel et al., 2008; Leskovec, 2021). We observe that Louvain obtains high-quality communities, with $3.0-30\%$ higher modularity than that obtained by LPA, but requires 2.314×2.3-14$\times$2.3 - 14 × longer to converge.
4. Approach
4.1. Optimizations for Louvain algorithm
We use a parallel implementation of the Louvain method to determine suitable parameter settings and optimize the original algorithm through experimentation with a variety of techniques. We use asynchronous version of Louvain, where threads work independently on different parts of the graph. This allows for faster convergence but can also lead to more variability in the final result (Blondel et al., 2008; Halappanavar et al., 2017). Further, we allocate a separate hashtable per thread to keep track of the delta-modularity of moving to each community linked to a vertex in the local-moving phase of the algorithm, and to keep track of the total edge weight from one super-vertex to the other super-vertices in the aggregation phase of the algorithm.
Our optimizations include using OpenMP’s dynamic loop schedule, limiting the number of iterations per pass to $20$, using a tolerance drop rate of $10$, setting an initial tolerance of $0.01$, using an aggregation tolerance of $0.8$, employing vertex pruning, making use of parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for finding community vertices and for storing the super-vertex graph during the aggregation phase, and using fast collision-free per-thread hashtables which are well separated in their memory addresses (Far-KV) for the local-moving and aggregation phases of the algorithm. Details on each of the optimizations is given below.
For each optimization, we test a number of relevant alternatives, and show the relative time and the relative modularity of communities obtained by each alternative in Figure 3. This result is obtained by running the tests on each graph in the dataset (see Table 1), $5$ times on each graph to reduce the impact of noise, taking their geometric mean and arithmetic mean for the runtime and modularity respectively, and representing them as a ratio within each optimization category.
4.1.1. Adjusting OpenMP loop schedule
We attempt static, dynamic, guided, and auto loop scheduling approaches of OpenMP (each with a chunk size of $2048$) to parallelize the local-moving and aggregation phases of the Louvain algorithm. Results indicate that the scheduling behavior can have small impact on the quality of obtained communities. We consider OpenMP’s dynamic loop schedule to be the best choice, as it helps achieve better load balancing when the degree distribution of vertices is non-uniform, and offers a $7\%$ reduction in runtime with respect to OpenMP’s auto loop schedule, with only a $0.4\%$ reduction in the modularity of communities obtained (which is likely to be just noise).
4.1.2. Limiting the number of iterations per pass
Restricting the number of iterations of the local-moving phase ensures its termination within a reasonable number of iterations, which helps minimize runtime. This can be important since the local-moving phase performed in the first pass is the most expensive step of the algorithm. However, choosing too small a limit may worsen convergence rate. Our results indicate that limiting the maximum number of iterations to $20$ allows for $13\%$ faster convergence, when compared to a maximum iterations of $100$.
4.1.3. Adjusting tolerance drop rate (threshold scaling)
Tolerance is used to detect convergence in the local-moving phase of the Louvain algorithm, i.e., when the total delta-modularity in a given iteration is below or equal to the specified tolerance, the local-moving phase is considered to have converged. Instead of using a fixed tolerance across all passes of the Louvain algorithm, we can start with an initial high tolerance and then gradually reduce it. This is known as threshold scaling (Lu et al., 2015; Naim et al., 2017; Halappanavar et al., 2017), and it helps minimize runtime as the first pass of the algorithm (which is usually the most expensive). Based on our findings, a tolerance drop rate of $10$ yields $4\%$ faster convergence, with respect to a tolerance drop rate of $1$ (threshold scaling disabled), with no reduction in quality of communities obtained.
4.1.4. Adjusting initial tolerance
Starting with a smaller initial tolerance allows the algorithm to explore broader possibilities for community assignments in the early stage, but comes at the cost of increased runtime. We find an initial tolerance of $0.01$ provides a $14\%$ reduction in runtime of the algorithm with no reduction in the quality of identified communities, when compared to an initial tolerance of $10^{-6}$.
4.1.5. Adjusting aggregation tolerance
The aggregation tolerance determines the point at which communities are considered to have converged based on the number of community merges. In other words, if too few communities merged this pass we should stop here, i.e., if $|V_{aggregated}|/|V|\geq$ aggregation tolerance, we consider the algorithm to have converged. Adjusting aggregation tolerance allows the algorithm to stop earlier when further merges have minimal impact on the final result. According to our observations, an aggregation tolerance of $0.8$ appears to be the best choice, as it presents a $14\%$ reduction in runtime, when compared to the aggregation tolerance being disabled ($1$), while identifying final communities of equivalent quality.
4.1.6. Vertex pruning
Vertex pruning is a technique that is used to minimize unnecessary computation. Here, when a vertex changes its community, its marks its neighbors to be processed. Once a vertex has been processed, it is marked as not to be processed. However, it comes with an added overhead of marking an unmarking of vertices. Based on our results, vertex pruning justifies this overhead, and should be enabled for $11\%$ improvement in performance. An illustration of vertex pruning optimization is shown in Figure 1.
Figure 1. Illustration of vertex pruning optimization: After processing vertex $1$, it’s unmarked. If vertex $1$ changes its community, its neighbors are marked for processing. Community membership of each vertex is depicted by border color, and marked vertices are highlighted in yellow.
4.1.7. Finding community vertices for aggregation phase
In the aggregation phase of the Louvain algorithm, the communities obtained in the previous local-moving phase of the algorithm are combined into super-vertices in the aggregated graph, with the edges between two super-vertices being equal to the total weight of edges between the respective communities. This requires one to obtain the list of vertices belonging to each community, instead of the mapping of community membership of each vertex that we have after the local-moving phase ends. A straight-forward implementation of this would make use of two-dimensional arrays for storing vertices belonging to each community, with the index in the first dimension representing the community id $c$, and the index in the second dimension pointing to the $n^{th}$ vertex in the given community $c$. However, this requires memory allocation during the algorithm, which is expensive. Employing a parallel prefix sum technique along with a preallocated Compressed Sparse Row (CSR) data structure eliminates repeated memory allocation and deallocation, enhancing performance. Indeed, out findings indicate that using parallel prefix sum along with a preallocated CSR is $\times2$ faster than using 2D arrays for aggregating vertices.
4.1.8. Storing aggregated communities (super-vertex graph)
After the list of vertices belonging to each community have been obtained, the communities need to be aggregated (or compressed) into super-vertices, such that edges between two super-vertices being equal to the total weight of edges between the respective communities. This is generally called the super-vertex graph, or the compressed graph. It is then used as an input to the local-moving phase of the next pass of the Louvain algorithm. A simple data structure to store the super-vertex graph in the adjacency list format would be a two-dimensional array. Again, this requires memory allocation during the algorithm, which is bad for performance. Utilizing two preallocated CSRs, one for the source graph and the other for the target graph (except the first pass, where the dynamic graph may be stored in any desired format suitable for dynamic batch updates), along with parallel prefix sum can help here. We observe that using parallel prefix sum along with preallocated CSRs for maintaining the super-vertex graph is again $\times2$ faster than using 2D arrays.
4.1.9. Hashtable design for local-moving/aggregation phases
One can use C++’s inbuilt maps as per-thread (independent) hashtables for the Louvain algorithm. But this has poor performance. So we use a key-list and a full-size values array (collision-free) to dramatically improve performance. However, if the memory addresses of the hashtables are nearby (Close-KV), even if each thread uses its own hashtable exclusively, performance is not as high. This is possibly due to false cache-sharing. Alternatively, if we ensure that the memory address of each hashtable are farther away (Far-KV), the performance improves. Our results indicate that Far-KV has the best performance and is $\times4$ faster than Map, and 1.3×1.3$\times$1.3 × faster than Close-KV hashtable implementations. An illustration of Far-KV hashtable is shown in Figure 2.
Figure 2. Illustration of collision-free per-thread hashtables that are well separated in their memory addresses (Far-KV), for two threads. Each hashtable consists of a keys vector, values vector (of size $|V|$), and a key count ($N_{0}$/$N_{1}$). The value associated with each key is stored/accumulated in the index pointed by the key. As the key count of each hashtable is updated independently, we allocate it separately on the heap to avoid false cache sharing. These are used in the local-moving and aggregation phases of our Louvain implementation.
Figure 3. Impact of various parameter controls and optimizations on the runtime and result quality (modularity) of the Louvain algorithm. We show the impact of each optimization upon the relative runtime on the left Y-axis, and upon the relative modularity on the right Y-axis.
4.2. Our optimized Louvain implementation
We now explain the implementation of GVE-Louvain in Algorithms 1, 2, and 3. A flow diagram illustrating the first pass of GVE-Louvain is shown in Figure 4.
4.2.1. Main step of GVE-Louvain
The main step of GVE-Louvain (louvian() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, and the aggregation phase. Here, the louvain() function takes the input graph $G$, and returns the community membership $C$ for each vertex. In line 15, we first initialize the community membership $C$ of each vertex in $G$, and perform passes of the Louvain algorithm, limited to $MAX\_PASSES$ (lines 16-26). In each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 17); and mark all vertices as unprocessed (line 18).
Next, in line 19, we perform the local-moving phase by calling louvainMove() (Algorithm 2), which optimizes community assignments. If the local-moving phase converged in a single iteration, global convergence is implied and we terminate the passes (line 20). Further, if the drop is community count $|\Gamma|$ is too small, we have reached the point of diminishing returns, and thus stop at the current pass (line 22).
In case convergence has not been achieved, we renumber communities (line 23), update top-level community memberships $C$ with dendrogram lookup (line 24), perform the aggregation phase by calling louvainAggregate() (Algorithm 3), and scale the convergence threshold for subsequent passes, i.e., perform threshold scaling (line 26). The next pass continues in line 16. At the end of all passes, we perform a final update of the top-level community memberships $C$ with dendrogram lookup (line 27), and return the top-level community membership $C$ of each vertex in $G$.
Algorithm 1 GVE-Louvain: Our parallel Louvain algorithm.
1:$G$: Input graph
2:$C$: Community membership of each vertex
3:$G^{\prime}$: Input/super-vertex graph
4:$C^{\prime}$: Community membership of each vertex in $G^{\prime}$
5:$K^{\prime}$: Total edge weight of each vertex
6:$\Sigma^{\prime}$: Total edge weight of each community
7:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
8:$H_{t}$: Collision-free per-thread hashtable
9:$l_{i}$: Number of iterations performed (per pass)
10:$l_{p}$: Number of passes performed
11:$\tau$: Per iteration tolerance
12:$\tau_{agg}$: Aggregation tolerance
13:
14:function louvain($G$)
15:     Vertex membership: $C\leftarrow[0..|V|)$ ; $G^{\prime}\leftarrow G$
[…]
17:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow vertexWeights(G^{\prime})$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$
18:         Mark all vertices in $G^{\prime}$ as unprocessed
19:         $l_{i}\leftarrow louvainMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime})$
20:         if $l_{i}\leq 1$ then break $\triangleright$ Globally converged?
21:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
22:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
23:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
24:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
25:         $G^{\prime}\leftarrow louvainAggregate(G^{\prime},C^{\prime})$
[…]
27:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
28:     return $C$
Algorithm 2 Local-moving phase of GVE-Louvain.
1:$G^{\prime}$: Input/super-vertex graph
2:$C^{\prime}$: Community membership of each vertex
3:$K^{\prime}$: Total edge weight of each vertex
4:$\Sigma^{\prime}$: Total edge weight of each community
5:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
6:$H_{t}$: Collision-free per-thread hashtable
7:$l_{i}$: Number of iterations performed
8:$\tau$: Per iteration tolerance
9:
10:function louvainMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime}$)
[…]
12:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
13:         for all unprocessed $i\in V^{\prime}$ in parallel do
14:              Mark $i$ as processed (prune)
15:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
[…]
17:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
18:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
19:              if $c^{*}=C^{\prime}[i]$ then continue
20:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
21:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
22:              Mark neighbors of $i$ as unprocessed
23:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
24:     return $l_{i}$
25:
26:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
27:     for all $(j,w)\in G^{\prime}.edges(i)$ do
28:         if $self$ or $i\neq j$ then $H_{t}\leftarrow H_{t}[C^{\prime}[j]]+w$
29:     return $H_{t}$
Algorithm 3 Aggregation phase of GVE-Louvain.
1:$G^{\prime}$: Input/super-vertex graph
2:$C^{\prime}$: Community membership of each vertex
3:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
4:$G^{\prime\prime}$: Super-vertex graph (weighted CSR)
[…]
6:$H_{t}$: Collision-free per-thread hashtable
7:
8:function louvainAggregate($G^{\prime},C^{\prime}$)
[…]
10:     $G^{\prime}_{C^{\prime}}.offsets\leftarrow countCommunityVertices(G^{\prime},C^{\prime})$
11:     $G^{\prime}_{C^{\prime}}.offsets\leftarrow exclusiveScan(G^{\prime}_{C^{\prime}}.offsets)$
12:     for all $i\in V^{\prime}$ in parallel do
13:         Add edge $(C^{\prime}[i],i)$ to CSR $G^{\prime}_{C^{\prime}}$ atomically
[…]
15:     $G^{\prime\prime}.offsets\leftarrow communityTotalDegree(G^{\prime},C^{\prime})$
16:     $G^{\prime\prime}.offsets\leftarrow exclusiveScan(G^{\prime\prime}.offsets)$
17:     $|\Gamma|\leftarrow$ Number of communities in $C^{\prime}$
18:     for all $c\in[0,|\Gamma|)$ in parallel do
19:         if degree of $c$ in $G^{\prime}_{C^{\prime}}=0$ then continue
20:         $H_{t}\leftarrow\{\}$
21:         for all $i\in G^{\prime}_{C^{\prime}}.edges(c)$ do
22:              $H_{t}\leftarrow scanCommunities(H,G^{\prime},C^{\prime},i,true)$
23:         for all $(d,w)\in H_{t}$ do
24:              Add edge $(c,d,w)$ to CSR $G^{\prime\prime}$ atomically
25:     return $G^{\prime\prime}$
4.2.2. Local-moving phase of GVE-Louvain
The pseuodocode for the local-moving phase of GVE-Louvain is presented in Algorithm 2, which iteratively moves vertices between communities to maximize modularity. Here, the louvainMove() function takes the current graph $G^{\prime}$, community membership $C^{\prime}$, total edge weight of each vertex $K^{\prime}$, and total edge weight of each community $\Sigma^{\prime}$ as input, and returns the number of iterations performed $l_{i}$.
Lines 11-23 represent the main loop of the local-moving phase. In line 12, we first initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 22-22, we iterate over unprocessed vertices in parallel. For each vertex $i$, we mark $i$ as processed - vertex pruning (line 14), scan communities connected to $i$ excluding self (line 15), determine the best community $*c*$ to move $i$ to (line 17), calculate the delta-modularity of moving $i$ to $*c*$ (line 18), and update the community membership (lines 20-21) of $i$, and mark its neighbors as unprocessed (line 22) if a better community was found. In line 23, we check if the local-moving phase has converged (locally). If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 24, we return the number of iterations performed $l_{i}$.
4.2.3. Aggregation phase of GVE-Louvain
Finally, the psuedocode for the aggregation phase is shown in Algorithm 3, which aggregates communities into super-vertices. Here, the louvainAggrega te() function takes the current graph $G^{\prime}$ and the community membership $C^{\prime}$ as input, and returns the super-vertex graph $G^{\prime\prime}$.
In lines 10-11, the offsets array for the community vertices CSR $G^{\prime}_{C^{\prime}}.offsets$ is obtained by first counting the number of vertices belonging to each community using countCommunity Vertices(), and then performing exclusive scan on the array. In lines 12-13, we iterate over all vertices in parallel and atomically populate vertices belonging to each community into the community graph CSR $G^{\prime}_{C^{\prime}}$. Next, we obtain the offsets array for the super-vertex graph CSR by over-estimating the degree of each super-vertex, i.e., by obtaining the total degree of each community with communityTotalDegree(), and then performing exclusive scan on the array (lines 15-16). This causes the super-vertex graph CSR to be holey, i.e., with gaps in between the edges and weights array of each super-vertex in the CSR. Then, in lines 18-24, we iterate over all communities $c\in[0,|\Gamma|)$ in parallel, and add all communities $d$ (with associated edge weight $w$) linked to each vertex $i$ belonging to community $c$ (with scanCommunities() defined in Algorithm 2) to the per-thread hashtable $H_{t}$. Once $H_{t}$ is populated with all communities (and associated weights) linked to community $c$, we add them as edges to super-vertex $c$ into the super-vertex graph $G^{\prime\prime}$ atomically. At the end, in line 25, we return the super-vertex graph $G^{\prime\prime}$.
Figure 4. A flow diagram illustrating the first pass of GVE-Louvain for a Weighted 2D-vector based or a Weighted CSR with degree based input graph. In the local-moving phase, vertex community memberships are updated until the total change in delta-modularity across all vertices reaches a specified threshold. Community memberships are then counted and renumbered. In the aggregation phase, community vertices in a CSR are first obtained. This is used to create the super-vertex graph stored in a Weighted Holey CSR with degree. In subsequent passes, the input is a Weighted Holey CSR with degree and initial community membership for super-vertices from the previous pass.
5. Evaluation
5.1. Experimental Setup
We use a server that has two $16$-core x86-based Intel Xeon Gold 6226R processors running at $2.90$ GHz. Each core has an L1 cache of $1$ MB, an L2 cache of $16$ MB, and a shared L3 cache of $22$ MB. The machine has $93.4$ GB of system memory and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 1 shows the graphs we use in our experiments. All of them are obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
Table 1. List of $13$ graphs obtained SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with ***). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities obtained using Louvain algorithm.
Graph $|V|$ $|E|$ $D_{avg}$ $|\Gamma|$
Web Graphs (LAW)
$*{}^{*}$ 7.41M 341M 41.0 4.24K
$*{}^{*}$ 18.5M 567M 16.1 42.8K
$*{}^{*}$ 22.7M 1.21B 28.2 3.66K
$*{}^{*}$ 39.5M 1.73B 23.7 20.8K
$*{}^{*}$ 118M 1.89B 8.6 2.76M
$*{}^{*}$ 41.3M 2.19B 27.9 5.28K
$*{}^{*}$ 50.6M 3.80B 38.5 3.47K
Social Networks (SNAP)
com-LiveJournal 4.00M 69.4M 17.4 2.54K
com-Orkut 3.07M 234M 76.2 29
Road Networks (DIMACS10)
asia_osm 12.0M 25.4M 2.1 2.38K
europe_osm 50.9M 108M 2.1 3.05K
Protein k-mer Graphs (GenBank)
kmer_A2a 171M 361M 2.1 21.2K
kmer_V1r 214M 465M 2.2 6.17K
(a) Runtime in seconds (logarithmic scale) with Vite, Grappolo, NetworKit, and GVE-Louvain
(b) Speedup of GVE-Louvain with respect to Vite, Grappolo, NetworKit.
(c) Modularity of communities obtained with Vite, Grappolo, NetworKit, and GVE-Louvain.
Figure 5. Runtime in seconds (logarithmic scale), speedup, and modularity of communities obtained with Vite, Grappolo, NetworKit, and GVE-Louvain for each graph in the dataset.
5.2. Comparing Performance of GVE-Louvain
We now compare the performance of GVE-Louvain with Vite (Louvain), Grappolo (Louvain), and NetworKit Louvain. For Vite, we convert the graph datasets to Vite’s binary graph format, run it on a single node with threshold cycling/scaling optimization, and measure the reported average total time. For Grappolo, we measure the run it on the same system, and measure the reported total time. For NetworKit, we use a Python script to invoke PLM (Parallel Louvain Method), and measure the total time reported with getTiming(). For each graph, we measure the runtime of each implementation five times, for averaging. We also record the modularity of communities obtained, as reported by each implementation.
Figure 5(a) shows the runtimes of Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, and GVE-Louvain on each graph in the dataset. Figure 5(b) shows the speedup of GVE-Louvain with respect to each implementation mentioned above. GVE-Louvain is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. On the sk-2005 graph, GVE-Louvain finds communities in $6.8$ seconds, and thus achieve a processing rate of $560$ million edges/s. Figure 5(c) shows the modularity of communities obtained with each implementation. GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite (especially on web graphs), and $0.6\%$ lower modularity than Grappolo and NetworKit (especially on social networks with poor clustering).
(a) Phase split
(b) Pass split
Figure 6. Phase split of GVE-Louvain shown on the left, and pass split shown on the right for each graph in the dataset.
Figure 7. Runtime $/|E|$ factor with GVE-Louvain for each graph in the dataset.
Figure 8. Overall speedup of GVE-Louvain, and its various phases (local-moving, aggregation, others), with increasing number of threads (in multiples of 2).
5.3. Analyzing Performance of GVE-Louvain
The phase-wise and pass-wise split of GVE-Louvain is shown in Figures 6(a) and 6(b) respectively. Figure 6(a) indicates that GVE-Louvain spends most of the runtime in the local-moving phase on web graphs, road networks, and protein k-mer graphs, while it devotes majority of the runtime in the aggregation phase on social networks. The pass-wise split (Figure 6(b)) indicates that the first pass dominates runtime on high-degree graphs (web graphs and social networks), while subsequent passes prevail in execution time on low-degree graphs (road networks and protein k-mer graphs).
On average, $49\%$ of GVE-Louvain’s runtime is spent in the local-moving phase, $35\%$ is spent in the aggregation phase, and $16\%$ is spent in other steps (initialization, renumbering communities, looking up dendrogram, and resetting communities) of the algorithm. Further, $67\%$ of the runtime is spent in the first pass of the algorithm, which is the most expensive pass due to the size of the original graph (later passes work on super-vertex graphs) (Wickramaarachchi et al., 2014).
We also observe that graphs with lower average degree (road networks and protein k-mer graphs) and graphs with poor community structure (such as com-LiveJournal and com-Orkut) have a larger $\text{runtime}/|E|$ factor, as shown in Figure 7.
5.4. Strong Scaling of GVE-Louvain
Finally, we measure the strong scaling performance of GVE-Louvain. To this end, we adjust the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the overall time taken for finding communities with GVE-Louvain, as well as its phase splits (local-moving, aggregation, others), five times for averaging. The results are shown in Figure 8. With 32 threads, GVE-Louvain obtains an average speedup of 10.4×10.4$\times$10.4 × compared to running with a single thread, i.e., its performance increases by 1.6×1.6$\times$1.6 × for every doubling of threads. Scaling is limited due to the various sequential steps/phases in the algorithm. At 64 threads, GVE-Louvain is impacted by NUMA effects, and offers speedup of only 11.4×11.4$\times$11.4 ×.
6. Conclusion
In conclusion, this study addresses the optimization of Louvain method, a high-quality community detection algorithm, in the shared memory setting. We consider 9 different optimizations, which significantly improve the performance of the local-moving and the aggregation phases of the algorithm. On a server with dual 16-core Intel Xeon Gold 6226R processors, comparison with competitive open source implementations (Vite and Grappolo) and packages (NetworKit) show that our optimized implementation of Louvain, which we term as GVE-Louvain, is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. In addition, GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite, and $0.6\%$ lower modularity than Grappolo and NetworKit. On a web graph with $3.8$ billion edges, GVE-Louvain identifies communities in $6.8$ seconds, and thus achieves a processing rate of $560$ million edges/s. In addition, GVE-Louvain achieves a strong scaling factor of 1.6×1.6$\times$1.6 × for every doubling of threads. Looking ahead, future work could focus of designing dynamic algorithms for Louvain to accommodate dynamic graphs which evolve over time. This would contribute to interactive updation of community memberships of vertices in real-world scenarios.
Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, and Souvik Karfa for their support.